OpenStudy (anonymous):

Choose the equation of the line passing through the point (-2, 4) and perpendicular to y = -1/2x + 4. y = 2x + 8 y = 2x y = -2x - 8 y = -2x

jimthompson5910 (jim_thompson5910):

Anything perpendicular to y = -1/2x+4 has a slope of 2 (flip the fraction and change the sign of -1/2 to get 2) y = mx+b y = 2x+b 4 = 2(-2)+b 4 = - 4+b 4+4 = b 8 = b b = 8 So the answer is y = 2x+8